OpenStudy (anonymous):

How much heat energy, in kilojoules, is required to convert 34.0g of ice at −18.0 ∘C to water at 25.0 ∘C ? Specific heat of ice: sice=2.09 J/(g⋅∘C) Specific heat of liquid water: swater=4.18 J/(g⋅∘C) Enthalpy of fusion: ΔHfus=334 J/g Enthalpy of vaporization: ΔHvap=2250 J/g

OpenStudy (aaronq):

you have to do this in 3 steps and all all the q's together. 

OpenStudy (aaronq):

and add all*

OpenStudy (anonymous):

ok thanks so much

OpenStudy (aaronq):

no problem, if you're stuck just ask again.

OpenStudy (anonymous):

so for the first part q=mCdT is it (34.0g * 2.09 J(g*C)*delta 43)?

OpenStudy (aaronq):

not quite, in the first part you're only going from -18 to 0 celsius

OpenStudy (anonymous):

ok so q= 1279.08 delta T would be +?

OpenStudy (aaronq):

\(T_f-T_i=0-(-18)=18\) so yes, positive.

OpenStudy (ookawaiioo):

aaronq you are a beast